# Identifying Dwarfs of MC Analog GalaxiEs (ID-MAGE): The Search for Satellites Around Low-Mass Hosts

### v3

@@ 3.3 Visual Inspection
Objects scoring below two-thirds agreement are considered false positives and are not presented here. Approximately 65% of the candidates that pass the initial screening also pass the Zooniverse visual inspection. The number of objects which pass each step are presented in Table 2. Overall, we identify one candidate satellite for approximately every objects detected by the algorithm, aligning with findings from Bennet et al. (2017) and similar searches (e.g., Vollmer et al. 2013; Merritt et al. 2014; van der Burg et al. 2016; Carlsten et al. 2022). Table 3 showcases our list of candidates, detailing each candidate’s proposed host, ratings, photometric properties, and whether it is confirmed as a satellite based on a distance/velocity measurement in the literature.
@@ 3.4 Survey Completeness
The completeness in absolute magnitude depends on distance and so varies with host (see Appendix B). At the median distance of our host galaxies (6.25 Mpc), our 90% completeness limit of 20.2 is M. We find that our survey is complete down to roughly M, assuming 0.25. Our recovery rates are somewhat lower than those reported by Bennet et al. (2017), and similar to the ones observed in the ELVES and ELVES-Dwarf surveys. This arises from the varying depths of the photometric data employed by these studies. Specifically, Bennet et al. (2017) utilized deep imaging from the Canada-France-Hawaii Telescope Legacy Survey (CFHTLS), and ELVES used a combination of DESI Legacy Survey dr-9 data and archival CFHT/MegaCam data. Our data is all from the Legacy Survey dr-10 fields, and so is closest to the ELVES survey in terms of photometric completeness.
In this paper, we focus on candidates that lie within the high completeness regime of our survey. Therefore, we only include candidates with and ″(as the completeness drops below 50% for objects smaller than 3.1″). Additionally, there are very few known galaxies with that have an less than 100 pc (e.g., Brasseur et al. 2011; Doliva-Dolinsky et al. 2023; Pace 2024). We consider candidates with pc significantly more likely to be background galaxies rather than satellites. Therefore, to remove likely interlopers from the sample, we impose an additional physical size criteria cut, assuming the candidate is at the distance of the presumed host. To accommodate uncertainties in our photometric measurements (see Section 4) and to avoid removing potential satellites, we include candidates with 100 pc. This cut removes 22 objects from the sample, all of which are around hosts within 5.5 Mpc. See Table 3 for the complete candidate sample organized by host and candidate quality.
3.5 Known Objects
As part of our search process, we also check for known candidate satellite galaxies that the search algorithm may have missed. We identify five galaxies that are not recovered by the algorithm due to their close proximity to bright objects, resulting in partial or full masking. Three of these galaxies have known velocities or distances that align with their assumed hosts, while the other two do not have distance/velocity measures. We include these five galaxies in Table 3. Additionally, our algorithm successfully identified many previously known galaxies, which are incorporated into the main sample upon passing visual inspection. Most of these recovered galaxies come from previous photometric searches for low surface brightness galaxies (LSBGs, e.g., SMUDGes; Zaritsky et al. 2023, MATLAS Duc et al. 2015; Habas et al. 2020) and lack known distances or velocities. The algorithm detects 89% of the SMUDGes LSBGs within the footprint of the survey with a g21.0. In total, we detect 355 candidates, with 264 newly identified candidate satellites and 16 galaxies with published distance/velocities, confirming them as satellites.
4 Properties of Candidate Satellites
@@
| MAGE0245-6030 | 02h45m12.55s | -60d30m58.32s | ESO115-G21 | 17.40.1 | 17.20.1 | 0.420.01 | -11.20.3 | 7.10.1 | 21.80.1 | 6.30.2 | 0.930.1 | H |
| MAGE0231-6205 | 02h31m08.18s | -62d05m16.44s | ESO115-G21 | 19.70.1 | 19.30.1 | 0.680.01 | -9.00.3 | 4.70.1 | 23.30.1 | 5.70.2 | 1.090.1 | F |
| MAGE0243-6111 | 02h43m20.33s | -61d11m38.04s | ESO115-G21 | 19.30.1 | 18.90.1 | 0.470.01 | -9.40.3 | 4.50.1 | 22.60.1 | 5.90.2 | 1.00.1 | F |
| MAGE0233-6125 | 02h33m04.68s | -61d25m34.68s | ESO115-G21 | 18.50.1 | 18.50.1 | 0.550.01 | -10.00.3 | 6.80.1 | 22.50.1 | 5.70.2 | 1.230.1 | H |
| MAGE0252-5526 | 02h52m13.66s | -55d26m58.92s | ESO154-G23 | 18.90.1 | 18.50.1 | 0.290.01 | -10.10.3 | 6.00.1 | 21.60.1 | 6.10.2 | 1.390.1 | F |
| MAGE0248-5403 | 02h48m38.18s | -54d03m44.64s | ESO154-G23 | 19.70.1 | 19.30.1 | 0.590.01 | -9.40.3 | 5.10.1 | 23.00.1 | 5.80.2 | 1.260.1 | F |
| MAGE0246-5456 | 02h46m45.10s | -54d56m01.68s | ESO154-G23 | 18.40.1 | 18.20.1 | 0.60.01 | -10.50.3 | 4.90.1 | 22.90.1 | 6.00.2 | 0.520.1 | H |
| MAGE0304-5335 | 03h04m01.27s | -53d35m05.64s | ESO154-G23 | 19.40.1 | 19.20.1 | 0.730.01 | -9.50.3 | 3.20.1 | 22.90.1 | 5.60.2 | 0.70.1 | F |
| MAGE0139-4510 | 01h39m51.86s | -45d10m27.84s | ESO245-G05 | 19.30.1 | 19.10.1 | 0.40.01 | -9.10.3 | 5.10.1 | 23.50.1 | 5.50.2 | 0.480.1 | H |
| MAGE0144-4314 | 01h44m08.14s | -43d14m01.32s | ESO245-G05 | 18.50.1 | 18.30.1 | 0.580.01 | -9.80.3 | 5.70.1 | 23.50.1 | 5.70.2 | 0.390.1 | H |
| MAGE0143-4324 | 01h43m29.23s | -43d24m34.92s | ESO245-G05 | 18.50.1 | 18.10.1 | 0.640.01 | -9.90.3 | 10.10.1 | 24.10.1 | 6.00.2 | 0.880.1 | H |
| MAGE0140-4221 | 01h40m57.41s | -42d21m05.40s | ESO245-G05 | 19.60.1 | 18.90.1 | 0.480.01 | -9.10.3 | 5.80.1 | 23.10.1 | 6.00.2 | 1.210.1 | F |
| MAGE0142-4316 | 01h42m08.59s | -43d16m43.68s | ESO245-G05 | 18.50.1 | 18.20.1 | 0.560.01 | -9.90.3 | 4.40.1 | 22.30.1 | 6.00.2 | 0.810.1 | F |
| MAGE0143-4413 | 01h43m51.86s | -44d13m48.36s | ESO245-G05 | 19.20.1 | 18.40.1 | 0.360.01 | -9.50.3 | 5.90.1 | 21.80.1 | 6.40.2 | 1.60.1 | F |
| MAGE0606-3246 | 06h05m59.78s | -32d46m38.64s | ESO364-G29 | 20.70.1 | 20.00.1 | 0.80.01 | -9.40.3 | 5.60.5 | 24.20.1 | 6.10.2 | 1.450.12 | H |
| MAGE0606-3252 | 06h06m40.37s | -32d52m26.04s | ESO364-G29 | 19.10.1 | 18.70.1 | 0.840.01 | -10.90.3 | 4.00.1 | 22.80.1 | 6.30.2 | 0.990.1 | F |
| MAGE0608-3247 | 06h08m44.93s | -32d47m04.56s | ESO364-G29 | 18.10.1 | 17.70.1 | 0.630.01 | -11.90.3 | 3.80.1 | 21.90.1 | 6.80.2 | 0.690.1 | F |
| MAGE0605-3249 | 06h05m47.54s | -32d49m45.12s | ESO364-G29 | 19.40.1 | 18.70.1 | 0.470.01 | -10.80.3 | 10.50.6 | 23.60.1 | 6.80.2 | 1.550.1 | H |
| MAGE0605-3220 | 06h05m46.30s | -32d20m48.84s | ESO364-G29 | 18.80.1 | 18.60.1 | 0.770.01 | -11.00.3 | 3.60.1 | 21.40.1 | 6.30.2 | 1.440.1 | F |
| MAGE0606-3257 | 06h06m16.39s | -32d57m51.48s | ESO364-G29 | 17.90.1 | 17.60.1 | 0.610.01 | -12.00.3 | 5.00.1 | 22.20.1 | 6.80.2 | 0.740.1 | H |
| MAGE0609-3225 | 06h09m09.91s | -32d25m22.08s | ESO364-G29 | 17.00.1 | 16.80.1 | 0.650.01 | -12.90.3 | 11.10.1 | 23.10.1 | 6.80.2 | 0.690.1 | H |
| MAGE0601-3306 | 06h01m52.34s | -33d05m59.64s | ESO364-G29 | 20.50.1 | 19.70.1 | 0.530.01 | -9.70.3 | 13.31.6 | 26.90.1 | 6.40.2 | 0.740.1 | F |
@@
Note. — Column 1: Galaxy Name. Column 2: the Right Ascension (J2000.0). Column 3: the Declination (J2000.0). Column 4: TRGB distances from Tully et al. (2009) in Mpc. Column 5: Apparent Ks-band magnitude. Column 6: Absolute Ks-band magnitude (Kourkchi & Tully, 2017). Column 7: Apparent B-band magnitude. Column 8: B-band extinctions. Column 9: Absolute B-band magnitude (Karachentsev et al., 2013). Column 10: The logarithm of the total stellar mass, which are determined using the Ks luminosity and and M⋆/L0.6, assuming M3.27 in the Vega system (Willmer, 2018). Column 11: Tidal Index from Karachentsev et al. (2013).
@@
We use GALFIT (Peng et al., 2010) to measure the structural parameters of our candidates. We implement a fitting approach akin to the first fitting stage in Khim et al. (2024), outlined as follows.
@@
To minimize the influence of our initial parameter selections, we repeat the fitting procedure six times using various initial estimations. We employ combinations of two different effective radii (30 and 50 pixels) and three surface brightness values at the effective radii (20, 25, 30 mag arcsec-2). We calculate the reduced chi-squared statistic, , within a circular region of radius 50 pixels (approximately 13.1″) centered on the image center and opt for the model with the lowest value. Occasionally, GALFIT yields a model fit with huge final parameter uncertainties. We only regard models where as meaningful. For the r-band, we use the effective radius, central position, Sérsic index, axis ratio, and position angle from the g-band fit and only fit for magnitude and background level. The GALFIT results for each candidate are compiled in Table 3. The errors reported here are based on the uncertainties returned by GALFIT. We adopt a minimum uncertainty of 0.1 when GALFIT reports a smaller value, as it significantly underestimates the true photometric errors (see Bennet et al. 2017; Zaritsky et al. 2023). We will determine and report the full errors when we finalize our sample after our follow-up campaign (see Section 6).
@@
To estimate the stellar mass of our galaxies, we use the color-corrected stellar mass calibration for low-mass galaxies published in de los Reyes et al. (2024):
5 The Satellite Candidates
Around our 36 host galaxies, we identify 355 satellite candidates in the full sample and 134 in the high-likelihood sample. This section provides an overview of our candidates and the satellite systems. We first summarize the number of satellites identified per host. In Section 5.1, we compare the photometric properties of our satellites with those from the ELVES survey and dwarf galaxies within 3 Mpc. In Section 5.2, we assess the possible contamination rate of our sample and use it to estimate the lower range of our satellite luminosity function. Finally, Section 5.3 discusses the upper range of our sample and the luminosity function derived from the high-likelihood candidates.
@@ 5 The Satellite Candidates
We identify 98 high-likelihood candidates around the 27 SMC-mass hosts. The richest SMC-mass satellite system is NGC 3738, which has 15 high-likelihood candidates. However, this high number appears to be driven by a significant number of background galaxies contaminating the system (see Sections 5.2). Following NGC 3738, the next richest SMC-mass system is IC 4182, with 8 high-likelihood candidates. The system with the fewest candidates is UGC 04426, which has no high-likelihood candidates. Including NGC 3738, the mean number of high-likelihood satellite candidates per SMC-mass host is 3.5, but excluding NGC 3738, the mean drops to 3 satellites per system.
5.1 Comparison with Known Local Volume Dwarf Galaxies
@@
Figure 6 compares the ID-MAGE high-likelihood sample to the ELVES sample and dwarf galaxies within 3 Mpc (Pace, 2024), demonstrating that their derived physical properties are generally similar. However, our sample contains fewer galaxies with 500 pc compared to both the ELVES sample and the Pace (2024) compilation. As shown in Figure 4, this difference is not due to the completeness of our survey, which is 6090% for the full sample and 50–70% for the high-likelihood sample in this region. It is predicted that satellites of LMC/SMC-mass hosts are less massive and fainter than the satellites of MW-mass hosts. Given the size-luminosity relation, it is not surprising that we do not identify many satellites with 500 pc. Additionally, as illustrated in Figure 6, most high-likelihood candidates with 500 pc are confirmed satellites, indicating that there are very few potential false positives in this region, unlike the less massive candidates where few have published distances/velocities.
5.2 Sample Contamination

### v5

@@ III.3 Visual Inspection
Objects scoring below two-thirds agreement are considered false positives and are not presented here. Approximately 65% of the candidates that pass the initial screening also pass the Zooniverse visual inspection. The number of objects which pass each step are presented in Table 2. Overall, we identify one candidate satellite for approximately every objects detected by the algorithm, aligning with findings from Bennet17 and similar searches (e.g., Vollmer13; 2Merrit14; vanderburg16; ELVES). Table 3 showcases our list of candidates, detailing each candidate’s proposed host, ratings, photometric properties, and whether it is confirmed as a satellite based on a distance/velocity measurement in the literature.
@@ III.4 Survey Completeness
The completeness in absolute magnitude depends on distance and so varies with host (see Appendix B). At the median distance of our host galaxies (6.25 Mpc), our 90% completeness limit of 20.2 is M. We find that our survey is complete down to roughly M, assuming 0.25. Our recovery rates are somewhat lower than those reported by Bennet17, and similar to the ones observed in the ELVES and ELVES-Dwarf surveys. This arises from the varying depths of the photometric data employed by these studies. Specifically, Bennet17 utilized deep imaging from the Canada-France-Hawaii Telescope Legacy Survey (CFHTLS), and ELVES used a combination of DESI Legacy Survey dr-9 data and archival CFHT/MegaCam data. Our data is all from the Legacy Survey dr-10 fields, and so is closest to the ELVES survey in terms of photometric completeness.
In this paper, we focus on candidates that lie within the high completeness regime of our survey. Therefore, we only include candidates with and ″(as the completeness drops below 50% for objects smaller than 3.1″). Additionally, there are very few known galaxies with that have an less than 100 pc (e.g., Brasseur; ADD23; Pace24). We consider candidates with pc significantly more likely to be background galaxies rather than satellites. Therefore, to remove likely interlopers from the sample, we impose an additional physical size criteria cut, assuming the candidate is at the distance of the presumed host. To accommodate uncertainties in our photometric measurements (see Section IV) and to avoid removing potential satellites, we include candidates with 100 pc. This cut removes 22 objects from the sample, all of which are around hosts within 5.5 Mpc. See Table 3 for the complete candidate sample organized by host and candidate quality.
III.5 Known Objects
As part of our search process, we also check for known candidate satellite galaxies that the search algorithm may have missed. We identify five galaxies that are not recovered by the algorithm due to their close proximity to bright objects, resulting in partial or full masking. Three of these galaxies have known velocities or distances that align with their assumed hosts, while the other two do not have distance/velocity measures. We include these five galaxies in Table 3. Additionally, our algorithm successfully identified many previously known galaxies, which are incorporated into the main sample upon passing visual inspection. Most of these recovered galaxies come from previous photometric searches for low surface brightness galaxies (LSBGs, e.g., SMUDGes; SMUDGESV, MATLAS Duc15; Habas20) and lack known distances or velocities. The algorithm detects 89% of the SMUDGes LSBGs within the footprint of the survey with a g21.0. In total, we detect 355 candidates, with 264 newly identified candidate satellites and 16 galaxies with published distance/velocities, confirming them as satellites.
IV Properties of Candidate Satellites
@@
| MAGE J0245-6030 | 02h45m12.55s | -60d30m58.32s | ESO115-G21 | 17.40.1 | 17.20.1 | 0.420.01 | -11.20.3 | 7.10.1 | 21.80.1 | 6.30.2 | 0.930.1 | H |
| MAGE J0231-6205 | 02h31m08.18s | -62d05m16.44s | ESO115-G21 | 19.70.1 | 19.30.1 | 0.680.01 | -9.00.3 | 4.70.1 | 23.30.1 | 5.70.2 | 1.090.1 | F |
| MAGE J0243-6111 | 02h43m20.33s | -61d11m38.04s | ESO115-G21 | 19.30.1 | 18.90.1 | 0.470.01 | -9.40.3 | 4.50.1 | 22.60.1 | 5.90.2 | 1.00.1 | F |
| MAGE J0233-6125 | 02h33m04.68s | -61d25m34.68s | ESO115-G21 | 18.50.1 | 18.50.1 | 0.550.01 | -10.00.3 | 6.80.1 | 22.50.1 | 5.70.2 | 1.230.1 | H |
| MAGE J0252-5526 | 02h52m13.66s | -55d26m58.92s | ESO154-G23 | 18.90.1 | 18.50.1 | 0.290.01 | -10.10.3 | 6.00.1 | 21.60.1 | 6.10.2 | 1.390.1 | F |
| MAGE J0248-5403 | 02h48m38.18s | -54d03m44.64s | ESO154-G23 | 19.70.1 | 19.30.1 | 0.590.01 | -9.40.3 | 5.10.1 | 23.00.1 | 5.80.2 | 1.260.1 | F |
| MAGE J0246-5456 | 02h46m45.10s | -54d56m01.68s | ESO154-G23 | 18.40.1 | 18.20.1 | 0.60.01 | -10.50.3 | 4.90.1 | 22.90.1 | 6.00.2 | 0.520.1 | H |
| MAGE J0304-5335 | 03h04m01.27s | -53d35m05.64s | ESO154-G23 | 19.40.1 | 19.20.1 | 0.730.01 | -9.50.3 | 3.20.1 | 22.90.1 | 5.60.2 | 0.70.1 | F |
| MAGE J0139-4510 | 01h39m51.86s | -45d10m27.84s | ESO245-G05 | 19.30.1 | 19.10.1 | 0.40.01 | -9.10.3 | 5.10.1 | 23.50.1 | 5.50.2 | 0.480.1 | H |
| MAGE J0144-4314 | 01h44m08.14s | -43d14m01.32s | ESO245-G05 | 18.50.1 | 18.30.1 | 0.580.01 | -9.80.3 | 5.70.1 | 23.50.1 | 5.70.2 | 0.390.1 | H |
| MAGE J0143-4324 | 01h43m29.23s | -43d24m34.92s | ESO245-G05 | 18.50.1 | 18.10.1 | 0.640.01 | -9.90.3 | 10.10.1 | 24.10.1 | 6.00.2 | 0.880.1 | H |
| MAGE J0140-4221 | 01h40m57.41s | -42d21m05.40s | ESO245-G05 | 19.60.1 | 18.90.1 | 0.480.01 | -9.10.3 | 5.80.1 | 23.10.1 | 6.00.2 | 1.210.1 | F |
| MAGE J0142-4316 | 01h42m08.59s | -43d16m43.68s | ESO245-G05 | 18.50.1 | 18.20.1 | 0.560.01 | -9.90.3 | 4.40.1 | 22.30.1 | 6.00.2 | 0.810.1 | F |
| MAGE J0143-4413 | 01h43m51.86s | -44d13m48.36s | ESO245-G05 | 19.20.1 | 18.40.1 | 0.360.01 | -9.50.3 | 5.90.1 | 21.80.1 | 6.40.2 | 1.60.1 | F |
| MAGE J0606-3246 | 06h05m59.78s | -32d46m38.64s | ESO364-G29 | 20.70.1 | 20.00.1 | 0.80.01 | -9.40.3 | 5.60.5 | 24.20.1 | 6.10.2 | 1.450.12 | H |
| MAGE J0606-3252 | 06h06m40.37s | -32d52m26.04s | ESO364-G29 | 19.10.1 | 18.70.1 | 0.840.01 | -10.90.3 | 4.00.1 | 22.80.1 | 6.30.2 | 0.990.1 | F |
| MAGE J0608-3247 | 06h08m44.93s | -32d47m04.56s | ESO364-G29 | 18.10.1 | 17.70.1 | 0.630.01 | -11.90.3 | 3.80.1 | 21.90.1 | 6.80.2 | 0.690.1 | F |
| MAGE J0605-3249 | 06h05m47.54s | -32d49m45.12s | ESO364-G29 | 19.40.1 | 18.70.1 | 0.470.01 | -10.80.3 | 10.50.6 | 23.60.1 | 6.80.2 | 1.550.1 | H |
| MAGE J0605-3220 | 06h05m46.30s | -32d20m48.84s | ESO364-G29 | 18.80.1 | 18.60.1 | 0.770.01 | -11.00.3 | 3.60.1 | 21.40.1 | 6.30.2 | 1.440.1 | F |
| MAGE J0606-3257 | 06h06m16.39s | -32d57m51.48s | ESO364-G29 | 17.90.1 | 17.60.1 | 0.610.01 | -12.00.3 | 5.00.1 | 22.20.1 | 6.80.2 | 0.740.1 | H |
| MAGE J0609-3225 | 06h09m09.91s | -32d25m22.08s | ESO364-G29 | 17.00.1 | 16.80.1 | 0.650.01 | -12.90.3 | 11.10.1 | 23.10.1 | 6.80.2 | 0.690.1 | H |
| MAGE J0601-3306 | 06h01m52.34s | -33d05m59.64s | ESO364-G29 | 20.50.1 | 19.70.1 | 0.530.01 | -9.70.3 | 13.31.6 | 26.90.1 | 6.40.2 | 0.740.1 | F |
@@
Note. — Column 1: Galaxy Name. Column 2: the Right Ascension (J2000.0). Column 3: the Declination (J2000.0). Column 4: TRGB distances from EDD in Mpc. Column 5: Apparent Ks-band magnitude. Column 6: Absolute Ks-band magnitude (Kourkchi17). Column 7: Apparent B-band magnitude. Column 8: B-band extinctions. Column 9: Absolute B-band magnitude (Karachen13). Column 10: The logarithm of the total stellar mass, which are determined using the Ks luminosity and and M⋆/L0.6, assuming M3.27 in the Vega system (Willmer18). Column 11: Tidal Index from Karachen13.
@@
We use GALFIT (Peng_2010) to measure the structural parameters of our candidates. We implement a fitting approach akin to the first fitting stage in khim+24, outlined as follows.
@@
To minimize the influence of our initial parameter selections, we repeat the fitting procedure six times using various initial estimations. We employ combinations of two different effective radii (30 and 50 pixels) and three surface brightness values at the effective radii (20, 25, 30 mag arcsec-2). We calculate the reduced chi-squared statistic, , within a circular region of radius 50 pixels (approximately 13.1″) centered on the image center and opt for the model with the lowest value. Occasionally, GALFIT yields a model fit with huge final parameter uncertainties. We only regard models where as meaningful. For the r-band, we use the effective radius, central position, Sérsic index, axis ratio, and position angle from the g-band fit and only fit for magnitude and background level. The GALFIT results for each candidate are compiled in Table 3. The errors reported here are based on the uncertainties returned by GALFIT. We adopt a minimum uncertainty of 0.1 when GALFIT reports a smaller value, as it significantly underestimates the true photometric errors (see Bennet17; SMUDGESV). We will determine and report the full errors when we finalize our sample after our follow-up campaign (see Section VI).
@@
To estimate the stellar mass of our galaxies, we use the color-corrected stellar mass calibration for low-mass galaxies published in dlReyes24:
V The Satellite Candidates
Around our 36 host galaxies, we identify 355 satellite candidates in the full sample and 134 in the high-likelihood sample. This section provides an overview of our candidates and the satellite systems. We first summarize the number of satellites identified per host. In Section V.1, we compare the photometric properties of our satellites with those from the ELVES survey and dwarf galaxies within 3 Mpc. In Section V.2, we assess the possible contamination rate of our sample and use it to estimate the lower range of our satellite luminosity function. Finally, Section V.3 discusses the upper range of our sample and the luminosity function derived from the high-likelihood candidates.
@@ V The Satellite Candidates
We identify 98 high-likelihood candidates around the 27 SMC-mass hosts. The richest SMC-mass satellite system is NGC 3738, which has 15 high-likelihood candidates. However, this high number appears to be driven by a significant number of background galaxies contaminating the system (see Sections V.2). Following NGC 3738, the next richest SMC-mass system is IC 4182, with 8 high-likelihood candidates. The system with the fewest candidates is UGC 04426, which has no high-likelihood candidates. Including NGC 3738, the mean number of high-likelihood satellite candidates per SMC-mass host is 3.5, but excluding NGC 3738, the mean drops to 3 satellites per system.
V.1 Comparison with Known Local Volume Dwarf Galaxies
@@
Figure 6 compares the ID-MAGE high-likelihood sample to the ELVES sample and dwarf galaxies within 3 Mpc (Pace24), demonstrating that their derived physical properties are generally similar. However, our sample contains fewer galaxies with 500 pc compared to both the ELVES sample and the Pace24 compilation. As shown in Figure 4, this difference is not due to the completeness of our survey, which is 6090% for the full sample and 50–70% for the high-likelihood sample in this region. It is predicted that satellites of LMC/SMC-mass hosts are less massive and fainter than the satellites of MW-mass hosts. Given the size-luminosity relation, it is not surprising that we do not identify many satellites with 500 pc. Additionally, as illustrated in Figure 6, most high-likelihood candidates with 500 pc are confirmed satellites, indicating that there are very few potential false positives in this region, unlike the less massive candidates where few have published distances/velocities.
V.2 Sample Contamination